#70 SEC.4.3
Directions: Find each of the following polynomial completely.
-3s^2-10s+8

Answer by checkley75(3666)

-3s^2-10s+8 you need to find thefactors for -3 & 8 that add up to -10
the factors for -3 are (-1*3) & (1*-3)
the factors for 8 is (1*8) & (-1*-8) & (2*4) & (-2*-4)
now find the pair that adds up to -10.
-1*1+3+8=-1+24=23 not a solution
-3*2+1*4=-6+4=-2 not a solution
1*2-3*4=2-12=-10 thus the factors are:
(-3s+2)(s+4)